OpenStudy (lilcathyx):

I am totally lost with this. can someone help ? & explain how to do it... :O A point Q on a segment with endpoints A (2, -1) and C (4, 2) partitions the segment in a 3:1 ratio. Find Q

OpenStudy (phi):

3:1 ratio means you divide the line between A and C into 4 equal parts, then start at point A and move 3 parts along it and put Q there.

OpenStudy (phi):

in turns out to be relatively easy. you can do the "x" distance and the "y" distance separately. first, find the distance from A to C in the x-direction. can you do that ?

OpenStudy (lilcathyx):

ill try :) hold on

OpenStudy (lilcathyx):

ok i used the distance formula \[d=\sqrt{(x2-x1)^2 + (y2-y1)^2}\] and i got sq root of 13 = 3.6

OpenStudy (lilcathyx):

that is not the one your asking me for right ?

OpenStudy (imstuck):

try this (hopefully it attached ok!)

OpenStudy (phi):

using the distance formula is a hard way to do this. just find the difference is the x values (as the first step) what do you get?

OpenStudy (imstuck):

It didn't; there's a formula that goes with it that explains how to do all of these with no problems with the ratio. It's really easy!!

OpenStudy (aum):

If \((x_1, y_1)\) and \((x_2, y_2)\) are the two points then, the x-distance = \(x_2-x_1\) the y-distance = \(y_2-y_1\)

OpenStudy (lilcathyx):

x- distance is 2 and y-distance is 3 ?

OpenStudy (aum):

correct.

OpenStudy (phi):

yes. x distance is 2. now divide that distance by 4 to get 2/ 4 = 1/ 2 move 3 "parts" or 3* 1/ 2 = 1 1 /2 add 1 1 /2 to the x value of point A that gives you the X value of C

OpenStudy (phi):

do the same for y y distance is 3 divide by 4 to get 3 /4 move 3 of these, i.e. 3 * 3 /4 = 9/ 4 add 9/ 4 to the y value of A to get the y value of C

OpenStudy (aum):

that gives you the x value of *Q* that gives you the y value of *Q*

OpenStudy (lilcathyx):

so i have to do this : since the x-distance is 2 i divide it by 4 2 /4 = 0.5 3 x 0.5 = 1.5 & then im lost -.-

OpenStudy (aum):

Add that 1.5 to the x-value of point A to get the x-value of point Q.

OpenStudy (lilcathyx):

1.5 + 2 = 3.5 ?

OpenStudy (aum):

Yes. Do the same for y.

OpenStudy (lilcathyx):

so i have to do this : since the x-distance is 2 i divide it by 4 2 /4 = 0.5 3 x 0.5 = 1.5 1.5 + 2= 3.5 y-distance is 3 3/4 =0.75 3 x 0.75=2.25 then 2.25 + 3 = 5.25 ?

OpenStudy (aum):

y-distance is 3. divide by 4 multiply by 3

OpenStudy (lilcathyx):

3/4 = 0.75 3 x 0.75 = 2.25 ?

OpenStudy (aum):

yes 2.25 + (-1) = 1.25

OpenStudy (phi):

***then 2.25 + 3 = 5.25 ?*** where did you get the 3. You add the 2.25 to the y-value of point A

OpenStudy (lilcathyx):

i got the 3 from the y- distance. 2.25 + (-1) = 1.25 oohhhh

OpenStudy (aum):



OpenStudy (lilcathyx):

just to make sure 2 /4 = 0.5 3 x 0.5 = 1.5 1.5 + 2= 3.5 y-distance is 3 3/4 =0.75 3 x 0.75=2.25 then 2.25 + (-1) = 1.25 ?

OpenStudy (aum):

Correct. The coordinates of Q are: (3.5, 1.25) or (7/2, 5/4)

OpenStudy (lilcathyx):

Cool! Thanks for helping ! & explaining ^.^ i understood everything

OpenStudy (aum):

You are welcome.